Assuming that the frequency v of vibrating string may depend upon (1) applied load f(2) length (1) of string and (3) was per unit length of string (4( prove that u k1|v✓f\u where u=m\1
1
Expert's answer
2020-12-24T11:38:20-0500

"\nu \propto F^al^bm^c"


"\dim F=[MLT^{-2}]"

"\dim l=[L]"

"\dim m=[ML^{-1}]"

"\dim \nu=[T^{-1}]"


"[M^0L^0T^{-1}]=[MLT^{-2}]^a[L]^b[ML^{-1}]^c=M^{a+c}L^{a+b-c}T^{-2c}\to"


"a+c=0"

"a+b-c=0"

"-2a=-1"


So, we have


"a=\frac{1}{2}"

"b=-1"

"c=-\frac{1}{2}"


"\nu \propto F^al^bm^c=\frac{1}{l}\sqrt{\frac{F}{m}}" . Answer